OpenStudy (caishaax3):

What is the simplified form of the fifth root of x to the fourth power times the fifth root of x to the fourth power?

OpenStudy (caishaax3):

\[\sqrt[5]{x^4?} \times \sqrt[5]{?x^4}\] is the problem

OpenStudy (caishaax3):

A.)\[\sqrt{x}\] B.)\[x \frac{ 1 }{ 3 }\] C.)\[x \sqrt[5]{x^3?}\] D.)\[x \frac{ 16 }{ 25 }\]

OpenStudy (zzr0ck3r):

What are ?

OpenStudy (jdoe0001):

\(\large \bf \sqrt[5]{x^4}\times \sqrt[5]{x^4}\implies \sqrt[5]{x^4\times x^4}\implies ?\)

OpenStudy (caishaax3):

i got \[x \frac{ 16 }{ 25 }\]

OpenStudy (caishaax3):

@jdoe0001

OpenStudy (jdoe0001):

hmmm 16/25? if you mean \(\large x^{\frac{16}{25}}\implies \sqrt[25]{x^{16}}\) and... dunno where the 16 and the 25 came from though, thus, no dice on that one

OpenStudy (jdoe0001):

so... just simplify the radicand at \(\large \sqrt[5]{x^4\times x^4}\) see what that gives you